Solve the partial differential equation



Px+q=p^2

1
Expert's answer
2021-11-17T13:16:30-0500

Its subsidiary equations are given by


"\dfrac{dx}{-f_p}=\dfrac{dy}{-f_q}=\dfrac{dz}{-pf_p-qf_q}=\dfrac{dp}{f_q+pf_z}"

"=\dfrac{dq}{f_y+qf_z}=\dfrac{d\phi}{0}"


"\dfrac{dx}{2p-x}=\dfrac{dy}{-1}=\dfrac{dz}{2p^2-xp-q}=\dfrac{dp}{p+0}"

"=\dfrac{dq}{0}=\dfrac{d\phi}{0}"

Taking "dq = 0 \u21d2 q = c (constant)"

"p^2-xp-q=0" becomes "p^2-xp-c=0"


"p=\dfrac{x\pm\sqrt{x^2+4c}}{2}"

Thus


"dz=pdx+qdy"

"dz=\dfrac{x\pm\sqrt{x^2+4c}}{2}dx+cdy"

Integrate


"z=\int(\dfrac{x\pm\sqrt{x^2+4c}}{2})dx+c\int dy+C_1"

"\int \dfrac{\sqrt{x^2+4c}}{2}dx=c\ln(|\sqrt{x^2+4c}+x|)+\dfrac{x}{4}\sqrt{x^2+4c}"

"z=\dfrac{x^2}{4}\pm(c\ln(|\sqrt{x^2+4c}+x|)+\dfrac{x}{4}\sqrt{x^2+4c})"

"+cy+C_2"